------------------------------------------------------------ Date: Sun Jan 11 08:03:02 2004, Shift: 00:00-08:00 Supervisor: sjs, with: Log book#15 pages 29 - 29 Shift Summary: ------------------------------------------------------------ 1. Two good stores overnight. Things ran smoothly. 2. Run 9541(FS@4/A1219;MRS@90/A1050-463K trig6) Run 9542(FS@4/A1723;MRS@90/A1050-24K trig6) Run 9547(FS@4/A1723;MRS@90/A1050-380K trig6) Run 9458(FS@4/A1723;MRS@90/A1050-329K trig6) Run 9549(FS@4/A1723;MRS@90/A1050-282K trig6/just finished) Pedestal runs 9543,9544 3. No new problems to report. Had to restart IOSERVER once, but daq recovered without further problem. At beginning of shift changed polarity of D5 from B to A. We do not have much of a run plan for the MRS, other than to let it sit. However, talking with CEJ, who is anxious for higher field running, it was decided to start on A after having worked on the B1050 at 90deg setting for several days. ------------------------------------------------------------ _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Sun Jan 11 08:03:19 2004
This archive was generated by hypermail 2.1.8 : Sun Jan 11 2004 - 08:03:37 EST